OpenStudy (anonymous):

Lt (-1)^n does not exist but what is the n->infty case with Lt (-1)^(n+1) n->infty

OpenStudy (jamesj):

Exactly the same.

OpenStudy (jamesj):

...because both alternate without end +1, -1, +1, -1, ...

OpenStudy (anonymous):

but how can u say the ans is zero in my previous question

OpenStudy (jamesj):

because it was all over (2n-1) and the limit as n --> infty of 1/(2n-1) is zero

OpenStudy (jamesj):

You should write out the first twenty terms or so of your series a_n = ( (-1)^n + 1 ) / ( 2n - 1 ) so you get a sense of of it. I think it's intuitively obvious that this does have limit zero. a1 = 0 a2 = 2/3 a3 = 0 a4 = 2/7 a5 = 0 a6 = 2/11 a7 = 0 a8 = 2/15 a9 = 0 a10 = 2/19 ....